OpenStudy (anonymous):

y = log x If y = 10, then what is x?

OpenStudy (ybarrap):

$$ \tt \Large{ y=\log x\\ 10=\log x\\ 10^{10}=10^{\log x}=x\\ So,~x=10^{10} } $$ Does this make sense?

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

x = 1

OpenStudy (ybarrap):

Log(1) = 0 http://www.wolframalpha.com/input/?i=log10%2810%5E10%29

OpenStudy (anonymous):

y = log x 10 = log x 10 = 10^x x = 1

OpenStudy (ybarrap):

Log answers the question, how many zeros are in 10^10 if y = 10. The answer is 10. So x = 10^10. 10^x is not log x if 10^x = log x then 10^(10^x) = 10 ^(log x)=x, which is not true

OpenStudy (ybarrap):

log(10)=1 log(100)=2 log(1000)=3 ... ... log(10^1)=1 log(10^2)=2 log(10^3)=3 ... log(10^x)=x

OpenStudy (ybarrap):

Do you see that? Punch in those numbers in a calculator and it will confirm (use base 10). This also means log(1)=log(10^0)=0